OpenStudy (anonymous):

Find the slope for 4.8x- 1.2y= 36 my answer is y=-4x+36 Is this correct? Thank you.

OpenStudy (bakonloverk):

@New_ You're answer is slightly off. 4.8x-1.2y=36 -4.8 -4.8 -1.2y=-4.8x+36 -4.8/-1.2 36/-1.2 so your answer would be y=4x-30

OpenStudy (anonymous):

You're being asked for the slope, not the slope intercept form of the equation. @bakonloverk is right about that form but the answer is slope=4